OpenStudy (anonymous):

Patrick is using a weed killer to treat his circular lawn which a diameter of 12 metres. The weed killer instructions say to use 20 ml for every square metre of lawn to be treated. How much weed killer does he need to use? Give your answer to the nearest ml.

OpenStudy (anonymous):

area = pi*r^2 area = pi*6^2 area=36pi m^2 weed killer = (20 ml )(36 pi m^2)

OpenStudy (anonymous):

720 ml

OpenStudy (anonymous):

oops forgot the pi

OpenStudy (anonymous):

20*36*pi = answer in ml

OpenStudy (anonymous):

diameter = 12m ==> radius = 6m area of circle = πr^2 = 3.14 * (6m)^2 = 3.14*36m^2 = 113.04m^2 for 1 m^2 area wed killer required = 20 ml so for 113.04m^2 area, weed killer required = (20 * 113.04) ml = 2260.8 ml = 2261 ml

OpenStudy (anonymous):

technically 2262 if you use full pi

OpenStudy (anonymous):

and dont round till the end like you're supposed to